OpenStudy (anonymous):

Use the fundamental identities to verify the following: cos theta sec theta - cos^2 theta = sin^2 theta and 1- (sin^2 theta)/1- cos theta = -cos theta

OpenStudy (mertsj):

Why don't we use x instead of theta.

OpenStudy (mertsj):

\[\cos x \times \sec x-\cos ^2x=\sin ^2x\] \[\cos x \times\frac{1}{\cos x}-\cos ^2x=\sin ^2x\] \[1-\cos ^2x=\sin ^2x\] \[\sin ^2x=\sin ^2x\]